Calculate the area of the triangle through the points (5,3,2),(-2,7,-1) and (4,-2,6)


1
Expert's answer
2021-08-18T05:47:52-0400

Solution.

We can use Heron formula which states

The area of a triangle with sides a,b,c is equal to

"S=\sqrt{p(p-a)(p-b)(p-c)}"


 where "p=\frac{a+b+c}{2}."

Using the formula to find the distance between two points "A(x_1,x_2,x_3), B(y_1,y_2,y_3)"

which is

"AB=\sqrt{(x_1-y_1)^2+(x_2-y_2)^2+(x_3-y_3)^2}"


we can calculate the length of sides between the three points given

let say A(5,3,2), B(-2,7,-1) and C(4,-2,6).

After that, we substitute to Heron formula.

"AB=\\sqrt{(5+2)^2+(3-7)^2+(2+1)^2}=\\sqrt{49+16+9}=\n\\sqrt{74}.\\newline\nBC=\\sqrt{(-2-4)^2+(7+2)^2+(-1-6)^2}=\\sqrt{36+81+49}=\n\\sqrt{166}.\\newline\nAC=\\sqrt{(5-4)^2+(3+2)^2+(2-6)^2}=\\sqrt{1+25+16}=\n\\sqrt{42}.\\newline\np = \t\\frac{\na + b + c}{\n2}\n = \t\\frac{\n1}{\n2}\n(\t8.6\t + \t12.88\t + \t6.48\t) = \t13.98\\newline\nS=\\sqrt{(13.98)(13.98 - 8.6)(13.98 - 12.88)(13.98 - 6.48) }=24.9 \\text{square units}\n\\newline\n\\text{Answer. 24.9 square units.}"